# Identifying Proca-star mergers via consistent ultralight-boson mass estimates across gravitational-wave events

### v1

@@ Abstract
While black-hole and neutron-star mergers are the most plausible sources of current gravitational-wave observations, mergers of exotic compact objects may mimic these signals. Proca stars – Bose-Einstein condensates of complex vector ultralight bosons – have gained significant attention for their potential to replicate certain gravitational-wave events while yielding consistent estimates of the boson mass $\mu_{B}$ forming the stars. Using a mixture model within a Bayesian framework, we demonstrate that consistent boson-mass estimates across events can yield conclusive evidence for the existence for a number $n$ of Proca-star families characterized by respective boson masses $\mu_{B}^{i}$, even if no individual event can be conclusively identified as such. Our method provides posterior distributions for $n$ and $\mu_{B}^{i}$. Applying this framework to the high-mass events GW190521, GW190426_190642, GW200220_061928, we obtain a Bayes Factor ${\cal{B}}^{n=0}_{n=1}=2$ against the Proca-star hypothesis, primarily rooted in the limitation of current Proca-star merger simulations to intrinsically weak head-on cases. We discuss the use of priors that mitigate the impact of these limitations, which yields ${\cal{B}}^{n=1}_{n=0}=8$ in favour of a single Proca family hypothesis; and the inclusion of the gravitational-wave trigger S200114__\__020818. Finally, we show that conclusive evidence $\log{\cal{B}}^{n=1}_{n=0}\geq 5$ could be achieved after 5 to 9 observations of similar event sets, at the $90\%$ credible level. This method provides a new way to detect exotic compact objects, somewhat using gravitational-wave detectors as particle detectors.
@@ I Introduction
The gravitational-wave (GW) detectors, Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015), now joined by KAGRA Akutsu et al. (2020); Somiya (2012), have made the observation of compact binary mergers almost routine. During their first three observing runs, these detectors have reported $\mathcal{O}(100)$ such observations that have provided us with unprecedented knowledge on how black holes (BHs) and neutron stars (NSs) form and populate our Universe Abbott et al. (2018, 2021a, 2021b). Moreover, these observations have enabled the first tests of General Relativity in the strong-field regime Abbott et al. (2021c, d) and qualitatively new studies of the Universe at a large scale Abbott et al. (2016, 2019).
@@
While BHs and NSs stand as the most plausible and anticipated GW sources, there exist further theoretical proposals of exotic compact objects (ECOs) that can mimic their GW emission Bezares and Sanchis-Gual (2024). A particularly appealing class of ECOs is that of boson stars Ruffini and Bonazzola (1969); Kaup (1968), which are among the simplest and most dynamically robust exotic compact objects one can consider. These consist on Bose-Einstein condensates of ultralight bosons that can clump together to form macroscopic objects of astrophysical size. Among boson stars, Proca-stars Brito et al. (2016) formed by complex-vector bosonic fields have recently arisen as extremely appealing candidates Brito et al. (2016) due to their particular properties. While scalar bosonic fields can form boson stars, these are unstable against non-axisymmetric perturbations and cannot therefore be stable if they have spin, unless self-interactions are considered Siemonsen and East (2021). Spinning Proca stars, in contrast, are stable and can therefore survive long enough to undergo mergers that can produce GW emission detectable by GW detectors Sanchis-Gual et al. (2017, 2019). Furthermore, vector bosons are not only motivated in some extensions of the Standard Model of elementary particles but have also been long considered as plausible candidates to form part of what we know as Dark Matter Arvanitaki et al. (2010); Freitas et al. (2021).
@@
Several groups have succeeded to obtain numerical simulations of boson-star mergers, extracting the corresponding GW emission Sanchis-Gual et al. (2019, 2022); Evstafyeva et al. (2024). The latter can then be used to construct templates to search for these objects in GW data and in parameter inference tasks Bustillo et al. (2021); Calderón Bustillo et al. (2023); Bustillo et al. (2023); Luna et al. (2024); Evstafyeva et al. (2024). In particular, although still restricted to the head-on merger case, both numerical simulations of Proca-star mergers (PSMs) and continuous surrogate waveform models have recently been compared to sufficiently short GW signals detected by LIGO and Virgo Bustillo et al. (2021); Calderón Bustillo et al. (2023); Luna et al. (2024); Bustillo et al. (2023). These analyses have found that the signal GW190521 Abbott et al. (2020a) is indeed consistent with a Proca-star merger, with this scenario being weakly preferred over the “vanilla” BBH one. In addition, a mass $\mu_{B}\simeq 8.7\times 10^{13}$ eV was estimated for the complex ultralight boson building up the stars. A recent follow-up was performed on two extra “catalogued” GW events, namely GW190426__\__190642 Abbott et al. (2024), GW200220__\__061928 Abbott et al. (2023), which we will refer to as simply GW200220 and GW190426 Interestingly, while showing statistical preference for the BBH scenario, these two events yield a boson-mass $\mu_{B}$ consistent with that for GW190521 (see Fig. 1). Finally, a fourth GW trigger known as S200114__\__020818 Abbott et al. (2022) (which we will refer to as S200114) weakly favours the PSM hypothesis, albeit yielding a boson mass inconsistent with GW190521.
@@
Notably, all of these events display component black-hole masses inconsistent with an stellar-collapse origin, as these fall within the so-called Pair-instability Supernova (PISN) Gap Heger et al. (2003). In this context, exotic-compact objects arise as a possible alternative explanation to hierarchical formation channels Gerosa and Fishbach (2021) which, while more plausible than ECOs, require restrictive merger configurations that lead to small gravitational recoils that prevent remnant black holes from leaving their host environments Araújo-Álvarez et al. (2024); Mahapatra et al. (2021, 2024). Similarly, ECOs shall also offer alternative explanations Mourelle et al. (2024) to low-mass objects that are slightly above the expected neutron-star mass and below that of stellar-born black holes Abbott et al. (2020b); Abac et al. (2024); Mahapatra et al. (2025).
@@
Figure 1: Boson-mass estimates for the events we consider in this work. Posterior distributions for the boson mass for the four events we consider, obtained under the assupmtion that these are sourced by head-on Proca-star mergers. Details on the corresponding analysis can be found in Calderón Bustillo et al. (2023).
@@
Both the gain in sensitivity of GW detectors towards their fourth observing run and the future existence of next-generation detectors like Cosmic Explorer Reitze et al. (2019); Evans et al. (2021), Einstein Telescope Hild et al. (2011); Punturo et al. (2010); Abac et al. (2025) or LISA Amaro-Seoane et al. (2017) make it reasonable to anticipate an important increase in the number of observed events similar to those discussed above. This, together with the development of improved templates for PSMs, will provide a great arena to probe the existence of potential sub-populations of PSMs mixed with “vanilla” BBHs. The identification of elements from such Proca-star populations will involve, at least, the comparison of individual events to both PSM and BBH models through a Bayesian framework to perform model selection. This will return posterior distributions on the individual source parameters including those for boson mass $\mu_{B}$. Model selection between BBH and PSM models across many observations may conclude that no individual event yields decisive evidence favouring the PSM model. On the one hand, this is may be due to the fact that current GW searches precisely target BBH signals Usman et al. (2016); Messick et al. (2017); Chandra et al. (2022), so that any detection should be sufficiently similar to a BBH. On the other hand, it is known that in some cases BBHs and PSMs can emit extremely similar signals Bustillo et al. (2021); Calderón Bustillo et al. (2023), making their discrimination intrinsically challenging. In this work we will show, however, that consistent estimations of parameters such as the boson mass ($\mu_{B}$) across PSM candidates can be exploited to conclusively identify specific sub-populations of PSMs characterized by distinct boson masses. Finally, even in a hypothetical situation where several conclusive PSM candidates already exist, our formalism can be leveraged to decide the number of ultralight bosons responsible for such candidates.
@@
The rest of this work is organised as follows. In section I, we introduce our working tool: a mixture model considering a mixed observation set consisting of BBHs mixed with $n$ families of PSM mergers respectively characterised q boson mass $\mu_{B}^{i}$ – with $i\in\{1,n\}$ – and a parameter $\zeta_{i}$ denoting the fraction of observations belonging to such family. Second, we apply this formalism to two different sets of existing GW observations that respectively exclude and include the trigger S200114. Performing analyses under allowing for different numbers $N_{\rm bosons}\equiv n$ ultralight bosons, we obtain Bayesian evidences for each value of $n$ and posterior distributions for the parameters $\mu_{B}^{i}$, $\zeta_{i}$ conditional to $n$. We also discuss the impact of prior choices. Finally, in Section III, we showcase the power of this formalism to detect PSM sub-populations on mock data sets consisting on several instances of event sets similar to the ones analysed in Section II. We show that the observation of 5-9 such sets would allow to a conclusive identification of a Proca-star merger sub-population while no evidence would be found if mass consistencies among events are ignored.
@@ II.1 A Proca-star counting experiment
$p(\{d_{i}\}|\zeta)$ (1)
$=\prod_{i=1}^{i=N}\bigg{[}p(d_{i}|\text{PSM})\zeta+p(d_{i}|\text{BBH})(1-\zeta)\bigg{]}$
$\propto\prod_{i=1}^{i=N}\bigg{[}{\cal{B}}^{\text{PSM}}_{\text{BBH}}\zeta+(1-\zeta)\bigg{]}.$
Above, the terms $p(d_{i}|\text{PSM})$ and $p(d_{i}|\text{BBH})$ denote respectively the probability of the data $d_{i}$ of the $i-th$ GW event given the PSM and BBH models against the noise hypothesis – also known as Bayesian evidence – while the term ${\cal{B}}^{\text{PSM}}_{\text{BBH}}$ denotes their ratio. In general, given a signal model $A$ depending on parameters $\vec{\theta}$, the Bayesian evidence is defined as
[…]
with $p(\vec{\theta}|A)$ denoting the prior probability for the parameters $\vec{\theta}$ and $p(d|\vec{\theta})$ denoting their likelihood given the data $d$.
@@ II.2 Counting Proca-stars under a common boson mass
Under the BBH scenario, the parameters $\vec{\theta}$ denote the two component masses and three-dimensional spins in addition to the extrinsic source parameters: the two sets of angles respectively parametrising the source orientation and sky-location, the luminosity distance, the coalescence time and the signal polarization. For the case of Proca stars we work with head-on configurations where both spins are aligned between them and with the final spin, which removes four spin parameters. Moreover, the star spin is linked to the mass-normalised oscillation frequency $\omega/\mu_{B}$ which, combined with the mass of the star determine the ultralight boson mass $\mu_{B}$.
@@
If Proca stars exist, it seems plausible that these may come in the form of a discrete number $n$ of sub-families characterised by a common ultralight boson with mass $\mu_{B}^{i}$. Here, $i\in\{1,...,N_{\rm bosons}=n\}$ with $N_{\rm bosons}$ denoting the number of ultralight bosons sourcing each of the $n$ Proca-star families. To start with, we will discuss the $n=1$ case, considering a single PSM family characterised by a boson mass $\mu_{B}$. Such model can be implemented by simply requiring that all analysed events share a common boson mass, as opposed to freely estimating such value for each individual event. Such condition has an important consequence in terms of model selection: it reduces number of sampled boson-mass parameters from $N$ (the number of considered events) to $n=1$, reducing the complexity of the model and, therefore, the incurred Occam Penalty. This will cause an increase of its evidence with respect to the BBH model if all events point to the same boson mass. On the contrary, if the individual events point to different masses, the model will not be able to fit the data making the evidence for the PSM model decrease.
@@
The joint likelihood for the fraction $\zeta$ of PSMs in our dataset and the common boson mass $\mu$ is given by
$p(\{d_{i}\}|\zeta,\mu)$ (3)
$=\prod_{i=1}^{i=N}\bigg{[}p(d_{i}|\text{PSM}(\mu))\zeta+p(d_{i}|\text{BBH})(1-\zeta)\bigg{]}$
$\propto\prod_{i=1}^{i=N}\bigg{[}{\cal{B}}^{\text{PSM}(\mu)}_{\text{BBH}}\zeta+(1-\zeta)\bigg{]},$
@@
Above, the quantity $p(d_{i}|\text{PSM}(\mu_{B}^{j}))$ denotes the Bayesian evidence for the PSM model assuming a boson mass prior consisting on a delta placed at $\mu_{B}=\mu$. This can be simply obtained from existing analyses making use of wider priors through the Savage-Dickey density ratio as
$p(d_{i}|\text{PSM}(\mu))=p(d_{i}|\text{PSM})\frac{p(\mu|d)}{p(\mu)},$ (4)
@@
where $p(\mu)$ denotes the prior on the boson mass $\mu$ imposed in such existing analyses and $p(\mu|d)$ denotes the corresponding posterior probability.

### v2

@@ Abstract
[…]
@@ I Introduction
The gravitational-wave (GW) detectors, Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015), now joined by KAGRA Akutsu et al. (2020); Somiya (2012), have made the observation of compact binary mergers almost routine. During their first three observing runs, these detectors have reported $\mathcal{O}(100)$ such observations that have provided us with unprecedented knowledge on how black holes (BHs) and neutron stars (NSs) form and populate our Universe Abbott et al. (2018, 2021a, 2021b); Collaboration et al. (2025a); Abbott et al. (2021c); Collaboration et al. (2021, 2025b). Moreover, these observations have enabled the first tests of General Relativity in the strong-field regime Abbott et al. (2021d, e) and qualitatively new studies of the Universe at large scales Abbott et al. (2017); Abbott and others. (2021); Abbott et al. (2023a).
@@
While BHs and NSs stand as the most plausible and anticipated GW sources, there exist further theoretical proposals of exotic compact objects (ECOs) that can mimic their GW emission Bezares and Sanchis-Gual (2024); Pani et al. (2010); Cardoso et al. (2016); Toubiana et al. (2021); Brito et al. (2016); Sanchis-Gual et al. (2019); Evstafyeva et al. (2024); Asali et al. (2020); Calderón Bustillo et al. (2023). A particularly appealing class of ECOs is that of boson stars Ruffini and Bonazzola (1969); Kaup (1968), which are among the simplest and most dynamically robust ECOs one can consider. These consist on Bose-Einstein condensates of ultralight bosons that can clump together to form macroscopic objects of astrophysical size. Among boson stars, Proca-stars Brito et al. (2016) formed by complex-vector bosonic fields have recently arisen as extremely appealing candidates Brito et al. (2016) due to their particular properties. While scalar bosonic fields can also form boson stars, they are unstable against non-axisymmetric perturbations and cannot therefore be stable if they have spin, unless self-interactions are considered Siemonsen and East (2021). Spinning Proca stars, in contrast, are stable and can therefore survive long enough to undergo mergers that can produce GW emission detectable by GW detectors Sanchis-Gual et al. (2017, 2019). Furthermore, vector bosons are not only motivated in some extensions of the Standard Model of elementary particles but have also been long considered as plausible candidates to form part of what we know as Dark Matter Arvanitaki et al. (2010); Freitas et al. (2021).
@@
Several groups have succeeded to obtain numerical simulations of boson-star mergers, extracting the corresponding GW emission Sanchis-Gual et al. (2019, 2022); Evstafyeva et al. (2024). The latter can then be used to construct waveform templates to search for these objects in GW data and in parameter inference tasks Bustillo et al. (2021); Calderón Bustillo et al. (2023); Bustillo et al. (2023); Luna et al. (2024); Evstafyeva et al. (2024). In particular, although still restricted to the head-on merger case, both numerical simulations of Proca-star mergers (PSMs) and continuous surrogate waveform models have recently been compared to sufficiently short GW signals detected by LIGO and Virgo Bustillo et al. (2021); Calderón Bustillo et al. (2023); Luna et al. (2024); Bustillo et al. (2023). These analyses have found that the signal GW190521 Abbott et al. (2020a) is indeed consistent with a head-on Proca-star merger, with this scenario being weakly preferred over the “vanilla” BBH one. In addition, a mass $\mu_{B}\simeq 8.7\times 10^{-13}$ eV was estimated for the complex ultralight boson building up the stars. A recent follow-up was performed on two extra “catalogued” GW events, namely GW190426_\_190642 Abbott et al. (2024), GW200220_\_061928 Abbott et al. (2023b), which we will refer to as simply GW200220 and GW190426. Interestingly, while showing statistical preference for the BBH scenario, these two events yield a boson-mass $\mu_{B}$ consistent with that inferred for GW190521 (see Fig. 1). Finally, a fourth GW trigger known as S200114_\_020818 Abbott et al. (2022) (which we will refer to as S200114) weakly favours the PSM hypothesis, albeit yielding a boson mass inconsistent with GW190521.
@@
Notably, all of the above events display component black-hole masses inconsistent with a stellar-collapse origin, as these fall within the so-called Pair-instability Supernova (PISN) gap Heger et al. (2003). In this context, exotic-compact objects arise as a possible alternative explanation to hierarchical formation channels Gerosa and Fishbach (2021) which, while more plausible than ECOs, require restrictive merger configurations that lead to small gravitational recoils Gonzalez et al. (2007); Varma et al. (2022); Bustillo et al. (2022) that prevent remnant black holes from leaving their host environments Araújo-Álvarez et al. (2024); Mahapatra et al. (2021, 2024). Similarly, ECOs shall also offer alternative explanations Mourelle et al. (2024) to low-mass objects that are slightly above the expected neutron-star mass and below that of stellar-born black holes Abbott et al. (2020b); Abac et al. (2024); Mahapatra et al. (2025).
@@
Figure 1: Boson-mass estimates for the events we consider in this work. Posterior distributions for the boson mass for the four events we consider, obtained under the assumption that these are sourced by head-on Proca-star mergers. Details on the corresponding analysis can be found in Calderón Bustillo et al. (2023).
@@
Both the gain in sensitivity of GW detectors towards their fourth observing run and the future existence of next-generation detectors like Cosmic Explorer Reitze et al. (2019); Evans et al. (2021), Einstein Telescope Hild et al. (2011); Punturo et al. (2010); Abac et al. (2025) or LISA Amaro-Seoane et al. (2017) make it reasonable to anticipate an important increase in the number of observed events similar to those discussed above. This, together with the development of improved templates for PSMs, will provide a great arena to probe the existence of potential sub-populations of PSMs mixed with “vanilla” BBHs. The identification of elements from such Proca-star populations will involve, at least, the comparison of individual events to both PSM and BBH models through a Bayesian framework to perform model selection. This will return posterior distributions on the individual source parameters including those for boson mass $\mu_{B}$. Model selection between BBH and PSM models across many observations may conclude that no individual event yields decisive evidence favouring the PSM model. On the one hand, this is may be due to the fact that current GW searches precisely target BBH signals Usman et al. (2016); Messick et al. (2017); Chandra et al. (2022), so that any detected GW should be sufficiently similar to a BBH. On the other hand, it is known that in some cases BBHs and PSMs can emit extremely similar signals Bustillo et al. (2021); Calderón Bustillo et al. (2023), making their discrimination intrinsically challenging. In this work we will show, however, that consistent estimations of parameters such as the boson mass ($\mu_{B}$) across PSM candidates can be exploited to conclusively identify specific sub-populations of PSMs characterized by distinct boson masses. Finally, even in a hypothetical situation where several conclusive PSM candidates already exist, our formalism can be leveraged to decide the number of ultralight bosons responsible for such candidates.
@@
The rest of this work is organised as follows. In section I, we introduce our working tool: a mixture model considering a mixed observation set consisting of BBHs mixed with $n$ families of PSM mergers respectively characterised a boson mass $\mu_{B}^{i}$ – with $i\in\{1,n\}$ – and a parameter $\zeta_{i}$ denoting the fraction of observations belonging to such family. Second, we apply this formalism to two different sets of existing GW observations that respectively exclude and include the trigger S200114. Performing analyses allowing for different numbers $n$ of ultralight bosons, we obtain Bayesian evidences for each value of $n$ and posterior distributions for the parameters $\mu_{B}^{i}$, $\zeta_{i}$. We also discuss the impact of prior choices. Finally, in Section III, we showcase the power of this formalism to detect PSM sub-populations on mock data sets consisting on several instances of event sets similar to the ones analysed in Section II. We show that the observation of 5-9 such sets would enable a conclusive identification of a Proca-star merger sub-population while no evidence would be found if mass consistencies across events are ignored.
@@ II.1 A Proca-star counting experiment
Consider a dataset consisting on $N$ GW observations $d=\{d_{i}\}$ with $i\in\{1,\ldots,N\}$. Next, consider a mixture model where a fraction $\zeta$ of the events correspond to PSMs while a fraction $1-\zeta$ correspond to BBHs. The likelihood for the fraction parameter $\zeta$ can be expressed as Callister et al. (2022):
$p(\{d_{i}\}|\zeta)$ (1)
$=\prod_{i=1}^{N}\bigg[p(d_{i}|\text{PSM})\zeta+p(d_{i}|\text{BBH})(1-\zeta)\bigg]$
$\propto\prod_{i=1}^{N}\bigg[{\cal{B}}^{\text{PSM}}_{\text{BBH},i}\zeta+(1-\zeta)\bigg].$
Above, the terms $p(d_{i}|\text{PSM})$ and $p(d_{i}|\text{BBH})$ denote respectively the probability of the data $d_{i}$ of the ithi-th GW event given the PSM and BBH models against the noise hypothesis – also known as Bayesian evidence – while the term ${\cal{B}}^{\text{PSM}}_{\text{BBH},i}$ denotes their ratio. In general, given a signal model $A$ depending on parameters $\vec{\theta}$, the Bayesian evidence is defined as
@@
with $p(\vec{\theta})$ denoting the prior probability for the parameters $\vec{\theta}$ and $p(d|\vec{\theta})$ denoting their likelihood given the data $d$.
@@ II.2 Counting Proca-stars under a common boson mass
Under the BBH scenario, the parameters $\vec{\theta}$ denote the masses and three-dimensional spins of the two component BHs in addition to the extrinsic source parameters: the two sets of angles respectively parametrising the source orientation and sky-location, the luminosity distance, the coalescence time and the signal polarization. For the case of PSMs, we work with results obtained in Calderón Bustillo et al. (2023) under the restriction to head-on merger configurations where both spins are aligned between them and with the final spin. In addition, the star spin is linked to the mass-normalised oscillation frequency of the boson field $\omega/\mu_{B}$. Combined with the mass of the star, $\omega/\mu_{B}$ determines the ultralight boson mass $\mu_{B}$. We note that the restriction of our analysis to head-on, aligned-spin configurations is caused by existing limitations in the obtention of reliable initial data to be used in numerical simulations of more generic configurations. This has to main effects. First, it will naturally limit the ability of PSM simulations to fit existing GW data, leading to small evidences for such scenario. Second, it will also potentially lead to over-constrained posterior distributions on the corresponding boson mass.
@@
If Proca stars exist, it seems plausible that these may come in the form of a discrete number of sub-families, each characterised by a value of the boson mass Arvanitaki et al. (2010); Freitas et al. (2021). This is, each family would be sourced by a common ultralight boson with mass $\mu_{B}^{i}$ with $i\in\{1,...,n\}$ and $n$ denoting the number of permitted ultralight bosons. To start with, we will discuss the $n=1$ case, where we consider a single PSM family characterised by a boson mass $\mu_{B}$. Such model can be implemented by simply requiring that all analysed events share a common boson mass $\mu_{B}$, as opposed to freely estimating such parameter for each individual event. Such condition has an important consequence in terms of model selection: it reduces the number of sampled boson-mass parameters from $N$ (the number of considered events) to $n=1$, reducing the complexity of the model and, therefore, the incurred Occam Penalty. This will cause an increase of its evidence with respect to the BBH model if all events point to the same boson mass. On the contrary, if the individual events point to different masses, the model will not be able to fit the data, making the evidence for the PSM model decrease.
@@
The joint likelihood for the fraction $\zeta$ of PSMs in our dataset and the common boson mass $\mu_{B}$ is given by
$p(\{d_{i}\}|\zeta,\mu_{B})$ (3)
$=\prod_{i=1}^{N}\bigg[p(d_{i}|\text{PSM}(\mu_{B}))\zeta+p(d_{i}|\text{BBH})(1-\zeta)\bigg]$
$\propto\prod_{i=1}^{N}\bigg[{\cal{B}}^{\text{PSM}}_{\text{BBH},i}(\mu_{B})\zeta+(1-\zeta)\bigg],$
@@
Above, the quantity $p(d_{i}|\text{PSM}(\mu_{B}))$ denotes the Bayesian evidence for the PSM model assuming a prior probability for the boson mass defined by a delta placed at a fixed value of $\mu_{B}$. This can be simply obtained from existing analyses making use of wider priors on $\mu_{B}$ through the Savage-Dickey density ratio as
$p(d_{i}|\text{PSM}(\mu_{B}))=p(d_{i}|\text{PSM})\frac{p(\mu_{B}|d_{i})}{p(\mu_{B})},$ (4)
@@
where $p(\mu)$ denotes the prior probability on the boson mass $\mu_{B}$ imposed in such existing analyses and $p(\mu_{B}|d_{i})$ denotes the corresponding posterior probability.